OpenStudy (anonymous):

3 4 - + - 4 5 Write As A Mixed Number

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 } + \frac{ 4 }{ 5 }\]Find the greatest common denominator.

OpenStudy (bahrom7893):

(15+16)/20

OpenStudy (anonymous):

Then add the two fractions.

OpenStudy (anonymous):

What do you get?

OpenStudy (bahrom7893):

sorry the previous one was a typo.. but jazy got this!

OpenStudy (anonymous):

So It Would Be 7 - 20

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

Ok, 20 is the GCD. So \[\frac{ 3 }{ 4 } = \frac{ ? }{ 20 }\] What is 3 *

OpenStudy (anonymous):

4 * 5 = 20 so multiply 3 by 5 as well

OpenStudy (anonymous):

What is 3 * 5